What is the factored form of 4x^2-32x+164
Found 2 solutions by Alan3354, Onmyoji.S:
Answer by Alan3354(69443)

What is the factored form of 4x^2-32x+164
----------
There's an obvious factor of 4, then it's prime.

Answer by Onmyoji.S(32)

4x²-32x+164=0
4x²-32x=-164
4(x²-8x)=-164
x²-8x=-41 (-8/2)²=16
x²-8x+16=-41+16 completing the square
(x-4)(x-4)=-25
(x-4)²=-25 factored form
x=+√25+4 x=+5+4 x=9
x=-√25+4 x=-5+4 x=-1 (√25=5)
so x=9,-1